Day shift of 31 Dec 1999

Leo Bellantoni (bellanto@fnal.gov)
Fri, 31 Dec 1999 16:02:40 -0600 (CST)

Shift Summary for Day shift, 31/Dec/99

On Shift:
Leo Bellantoni

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
None

========================================================================
Runs ended this shift:
None

Tapes brought to FCC this shift: KQ1001, ...
None

========================================================================
Accesses during shift:
None

========================================================================
Overview of shift:
A.C. failure, BUSY/deadtime studies

========================================================================
Description of Shift:

A.C. failure in counting room at start of shift due to failure of south
side chiller. Duty mechanic fixed it quickly; following this, some of the
usual struggles to get the front end started.

Study of ADC busys for the crates with recently installed 11 bit ADCs for
use in KaMI. Rather tentative conclusion is that we will not hurt the
E799 live time unless the occupancy in these crates goes to over ~20
channels.

At end of shift, no triggers being processed on KTeV1, 2, 4;
drt_fresh_start in progress.

========================================================================
Follow up on Notes from Previous Shift

========================================================================
Items to Follow up on Next Shift

Keep cosmics going.

========================================================================
Quote of the day:
"Why yes, Cololnel Steve Austin - that's a very good question."